40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Mon Aug 3 13:10:46 2020, shruti, DailyProgress, Lab Work, Initial experiments and prep IMG_633BBA57553C-1.jpeg
    Reply  Tue Aug 4 14:08:44 2020, shruti, DailyProgress, Lab Work, Initial experiments and prep: Beam profile IMG_0268.jpgErf.pdfDiv.pdfbeamprofdata.zipIMG_0258.jpg
Message ID: 2548     Entry time: Mon Aug 3 13:10:46 2020     Reply to this: 2549
Author: shruti 
Type: DailyProgress 
Category: Lab Work 
Subject: Initial experiments and prep 

Notice of lab entry: I will be at the cryo this (03 Aug 20) afternoon

Lights fixed, safety, cleanliness:

- On entering the lab I noticed that the work with the light fixtures was completed. Unfortunately I could not avert it or get it all covered in time; I had assumed that I would be contacted beforehand but was not. But, by inspecting the table I do not think it looked any dustier than before. For any such activity in the future after we’ve cleaned the optics, I will remember to get it covered beforehand.

- The particle counter did not seem to be saving any data so I’m unsure what the effect of this activity was. The 1 micron particle count:
When I entered (1 pm): 80
(6 pm) : 70
This change might probably be entirely attributed to the air scrubber.

- When Aaron and I chatted with Calum today about COVID safety, Calum pointed out that turning on any HEPA filters in the lab would reduce the time between single-person occupancy to almost 0 for our room, as it would serve as an additional air scrubber. I tried locating a switch on the cryo cavity table for the HEPA fans but could not locate it today.

Laser and beam profile related:

- After reading through bits of the laser operating manual and turning on the east cavity laser, I turned on the TEC (thermo-electric cooler) box [TED 200 C] with neither the TEC nor the servo on. I used this just to read off the resistance values of the sensor. With the current drive off it was 11.2 k-ohms [22 C].

- I waited until the sensor readings stabilized to changing at the level of a few ohms around 8.5 k-ohms [28 C] in a minute and began measuring the beam profile. (See Attachment 1 for setup)
I could not mount properly the only translation stage I found since the holes on it were smaller than 8-32, and moreover, I could not locate anything suitable that fits into the 1/4" holes on the table with this.

- Nonetheless, when I took some data and plotted it, it looked like an error function so I took more data. Will post plots later after fitting and analyzing. I may have to repeat this again.

- The sensor for the power measurement was S122C (Germanium) with a range of 700-1800 nm and 40 mW. After setting the power meter to 1550 nm, the measured power of the entire cross-section was ~2.3 mW.

- Also in the image in Attachment 1, the two optics (HWP and mirror) near the fiber-free space coupler, were previously in the path of the laser and now moved to either side of the beam. No other optics were moved from the previous set-up.

Attachment 1: IMG_633BBA57553C-1.jpeg  1.248 MB  Uploaded Mon Aug 3 19:12:27 2020  | Hide | Hide all
IMG_633BBA57553C-1.jpeg
ELOG V3.1.3-